OpenStudy (anonymous):

(x^2+x)/(x^2-4)/(x^2-1)/(x^2+5x+6)

jhonyy9 (jhonyy9):

like a first step x^2 +x = x(x+1) x^2 -4 =(x-2)(x+2) x^2 -1 =(x-1)(x+1) x^2 +5x +6 =(x+3)(x+2) hope these will help you

OpenStudy (anonymous):

Yes! Thank you so much